OpenStudy (anonymous):

5(3y-2)=16y need help

OpenStudy (anonymous):

15y-10=16y ....-10=16y-15y......-10=y

OpenStudy (anonymous):

-10

OpenStudy (accessdenied):

Notes for this problem: Distributive Property: a(b + c) = ab + ac If you distribute the 5 into the expression 3y - 2, you may solve for y by isolating the y-terms on one side (subtracting y-terms from both sides to move them) and dividing by the coefficient of y (if necessary) in order to get y alone.